OpenStudy (gabylovesyou):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = 1/4x + 2. y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (anonymous):

take the negative reciprocal of the slope. sub the given point in and solve for your constant.

OpenStudy (gabylovesyou):

......

hartnn (hartnn):

first,can u find the slope of the given line y=1/4 x + 2?

OpenStudy (gabylovesyou):

i can find the slope if u give me 2 coordinates but not like that

hero (hero):

@hartnn, she doesn't have to find the slope of that line. It's already given.

hero (hero):

You're asking her the wrong question.

OpenStudy (gabylovesyou):

1/4

hero (hero):

Yes @Gabylovesyou

OpenStudy (gabylovesyou):

ok.. now what lol

hartnn (hartnn):

ok,i was trying to use the fact that slopes of perpendicular lines have product -1.....so that she gets slope of required line......

OpenStudy (anonymous):

@Hero , help me out with my question? vectors are a nightmare for me

OpenStudy (gabylovesyou):

@hartnn now what?? the slope is 1/4

hartnn (hartnn):

ok,since the product of slope of perpendicular lines is -1,can u find the slope of required line??

OpenStudy (gabylovesyou):

how??

hartnn (hartnn):

let the slope of your line be m since product is -1 m*(1/4)=-1 so m=-1*4 = -4 did u get this?

OpenStudy (gabylovesyou):

yessss

hartnn (hartnn):

good,now we can take the slope-point form of the equation y=mx+c u found m as -4,only remaining work is to find c to find c,we put the point (x,y)=(-1,-2) like this: -2=(-4)(-1)+c can u find c from here?

OpenStudy (gabylovesyou):

-6

OpenStudy (anonymous):

i think it's probably over 9000

hartnn (hartnn):

yes,it is -6 so your equation y=mx+c with m=-4,c=-6 becomes y=-4x-6 thats your 2nd option :)

OpenStudy (gabylovesyou):

thank you soo much!! :)

hartnn (hartnn):

did u understand all the steps clearly?? ask if u still have doubts..

OpenStudy (gabylovesyou):

yea i did thanks soo much

hartnn (hartnn):

welcome :)

OpenStudy (gabylovesyou):

i have to do another one of those :S

hartnn (hartnn):

try it,if u don't get it,we are here :)

OpenStudy (gabylovesyou):

ok thanks

OpenStudy (gabylovesyou):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -1/4x - 5.

OpenStudy (gabylovesyou):

y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (gabylovesyou):

is it A

hartnn (hartnn):

no,what did u find slope of 2nd line as?

OpenStudy (gabylovesyou):

-1/4 for the slope of y = -1/4x - 5.

hartnn (hartnn):

so now what will be the slope of the line perpendicular to it?...remember that product of slopes of perpendicular lines = -1 as already told u

OpenStudy (gabylovesyou):

mmmmmm what would be the equation?

hartnn (hartnn):

let the slope of your line be m since product is -1 m*(-1/4)=-1 so m=-1*-4 = 4

OpenStudy (gabylovesyou):

OMG i knew it it was 4!!!

hartnn (hartnn):

now put -1,7 in the equation y=4x+c to get c.....

OpenStudy (gabylovesyou):

7 = 4*-1 + C ??

OpenStudy (gabylovesyou):

C = 11 is the answer C

hartnn (hartnn):

very good :) so whats the final answer?

OpenStudy (gabylovesyou):

third option??

hartnn (hartnn):

yup :)